the acceleration of a certain marble is given by,


a=-3(velocity squared)


if the marble enters the fluid with a speed of 1.5m/s , how long will it take before marble's speed is reduced to half of its initial value?

Grade:

2 Answers

SAGAR SINGH - IIT DELHI
13 years ago

Dear student,

a=-3v2

dv/dt=-3v2

dv/v2=-3dt

-1/v=-3t

So u get a relation between velocity and time..

Please feel free to ask your queries here. We are all IITians and here to help you in your IIT JEE preparation.

All the best.

Win exciting gifts by answering the questions on Discussion Forum. So help discuss any query on askiitians forum and become an Elite Expert League askiitian.

Now you score 5+15 POINTS by uploading your Pic and Downloading the Askiitians Toolbar  respectively : Click here to download the toolbar..

 

Askiitians Expert

Sagar Singh

B.Tech, IIT Delhi

vikas askiitian expert
13 years ago

a = dv/dt

a = -3v2            (given)

dv/dt = -3v2

dv/v2 = -3dt

-1/v = -3t + c        .............1

at t=0 , v = 1.5

c = -2/3

now eq 1 becomes

-1/v = -3t + -2/3           ................2

finally velocity becomes half of original= v = 1.5/2

after putting this in eq 2 we get

t = 2/9 sec

t =